# AutoTimes: Autoregressive Time Series Forecasters via Large Language Models

### v2

@@ 4.1 Time Series Forecasting
For short-term forecasting, we follow the well-acknowledged TimesNet [41], which assesses the fundamental ability of forecasters in modeling temporal variations. For long-term forecasting, we establish a novel one-for-all benchmark: a single forecaster is trained on one dataset and subsequently utilized for all prediction lengths. We highlight that this approach evaluates the basic versatility as foundation models of time series, which aims to break the prevailing practice of extensive training across diverse real-world scenarios. To be specific, we evaluate all methods by rolling forecasting: a model is trained with predetermined input/output lengths, and the forecasted values are integrated as part of the input in subsequent iterations until reaching the desired forecast length. Therefore, the key to success in this task lies in mitigating multi-step error accumulation. Still, the conventional one-for-one approach that trains forecasters respectively on each length is also provided in Table 11.
@@ Results
By diving into the proposed one-for-all and the traditional one-for-one benchmarks in Table 3 and 11, it is notable that prevalent deep forecasters, such as Transformer-based forecasters and DLinear, can achieve competitive and even better results under rolling forecasting. Nevertheless, the performance of non-autoregressive LLM4TS methods can degenerate a lot without respective training. Therefore, it highlights our persistent utilization of autoregression and thorough leveraging of inherent token transitions of LLMs, thereby mitigating error accumulation during multi-step rolling forecasting.
@@ 4.3 In-context Forecasting
We adopt the aforementioned M4 $\to$ M3 scenario. Since the samples of the M3 dataset are univariate time series with different lengths, we always predict the last $F$ time points of each sample during inference. We set the lookback length $L=F$, and thus the length of time series prompts is $2F$. We set the number of prompts $m=1$. Therefore, we initially train an LLM-based forecaster on the source M4 dataset with the context length of $3F$. We adopt an intuitive strategy to select the prompt: uniformly adopting the first $2F$ time points of the time series as the corresponding prompt. Supposing the lookback series starts after time $t$, the model for in-context learning is formulated as:
$f:(\{x_{1:2F}\},x_{t+1:t+F},\mathbf{a}_{t+1:t+2F})\mapsto\hat{x}_{t+F+1,t+2F},$ (12)
To prevent data leakage of future information, too short samples are discarded to prevent the overlap between the prompt and the lookback window. Implementation details are provided in Appendix D.6.
@@ 4.4 Method Analysis
Table 6: Averaged long-term forecasting results of AutoTimes and AutoTimes equipped with LoRA.
Datasets
@@
AutoTimes
0.397 0.425 0.173 0.266 0.242 0.278 0.406 0.276 0.207 0.246
AutoTimes + LoRA
0.396 0.425 0.161 0.255 0.231 0.268 0.396 0.275 0.201 0.243
@@ Appendix A Dataset Descriptions
Table 7: Detailed dataset descriptions. Dim denotes the variate number. Dataset Size denotes the total number of time points in (Train, Validation, Test) split respectively. Forecast Length denotes the future time points to be predicted. Frequency denotes the sampling interval of time points.
@@ Appendix B Implementation Details
AutoTimes processes timestamps in textual form rather than numerical encoding, potentially enabling handling other textual data such as news or logs. We utilize LLM to obtain embedding for the special token <EOS> to capture embedding for the entire sentence. Pseudo-code for this process is depicted in Algorithm 1. It’s worth noting that in the context of multivariate time series forecasting, timestamps are shared across variates. Thus, timestamps can implicitly express relationships between variates even with channel independence. Further, assuming there are $C$ variates since the number of timestamps is $\frac{1}{C}$ of the total time point count, its embedding can be efficiently pre-computed by large language models.
After obtaining embedding for the timestamps, we repurpose LLM for time series forecasting using Algorithm 2. At this stage, only the parameters of $\operatorname{SegmentEmbedding}$ and $\operatorname{SegmentProjection}$ are updated, while the parameters of LLM remain entirely frozen. During inference, AutoTimes utilizes the last token generated as its prediction and then employs this output to create subsequent predictions autoregressively. This approach enables AutoTimes to predict sequences of variable lengths with just one model dynamically. Such capability is crucial in real-world application scenarios. The pseudo-code in Algorithm 3-4 illustrates this process.
@@
2:$\mathbf{s}_{i}=\{x_{(i-1)S+1},\dots,x_{iS}\}$triangleright Dimensionality of LLM hidden states $D$
@@
1:Input time series $\{x_{1},\dots,x_{N_{1}\times S}\}$; text embeddings $\{\mathbf{TE}_{1},\dots,\mathbf{TE}_{N_{2}}\}$triangleright Forecast token number $N_{2}-N_{1}$
@@ D.1 Time Series Forecasting
We compare the performance of AutoTimes with state-of-the-art LLM4TS methods and well-acknowledged deep forecasters. Table 10 shows detailed short-term forecast results on M4. Table 9 presents results of the one-for-all forecasting benchmark across ETTh1, ECL, Traffic, Weather, and Solar-Energy datasets. We evaluate all methods by rolling forecasting: each model is trained with input length $672$ and output length $96$, and the forecasted values are integrated as part of the input in the next iteration until reaching the desired forecast length.
@@
Method AutoTimes TimeLLM [15] UniTime [21] FPT [48] iTrans. [22] DLinear [44] PatchTST [26] TimesNet [41]
@@
Table 11: Long-term forecasting results of one-for-one: AutoTimes trains one LLM-based forecaster to handle all prediction lengths by autoregression, whereas other models are trained respectively on each prediction length. AutoTimes adapt LLMs with the context length $C=672$. The lookback length is set as $L=672$ in others. All results are averaged. Full results is provided in Table 13.
@@ D.2 Zero-shot Forecasting
Following the zero-shot forecasting of FPT [15], every experiment comprises the source and target datasets. We train a model on the source dataset and apply the model on the target dataset for predictions directly.
@@ D.3 Method Generality
Mainstream LLMs predominantly adopt the decoder-only architecture, and AutoTimes can utilize any decoder-only LLM. We conduct experiments on various types and sizes of LLMs, including GPT-2 [31], multiple sizes of OPT [45], and LLaMA [35]. Specific results are shown in Table 15, demonstrating a general trend where performance improves as the model size increases, consistent with the scaling law [16].
@@ D.4 Variable Lookback Length
In the conventional forecasting paradigm, deep forecasters are trained respectively on lookback/forecast lengths, limiting their applicability to a single lookback length. In contrast, LLMs have the versatility to handle various input lengths. This capability is derived from Rotary Position Embedding  [33] and the next toke prediction objective, where LLMs are trained with token-wise supervision in Equation 8, that is, the generated token at each position is supervised. While non-autoregressive LLM4TS methods are typically constrained to a fixed lookback setting, AutoTimes with a consistent training objective has the flexibility to deal with different lookback lengths. We present the results in Figure 7, where we adapt the LLM by AutoTimes with the context length of $C=672$, and evaluate the performance with different lookback lengths, which demonstrates the inherited versatility of LLM-based forecasters. Moreover, the performance is generally improving with increased available lookback observations, leading to an averaged $9.3\%$ promotion from $384$ to $672$. Instead, several works have observed that the performance of respectively trained deep forecasters does not necessarily improve with the increasing of lookback length [22, 26, 44].
@@ D.6 In-context Forecasting
For in-context forecasting, similar to zero-shot forecasting in Appendix D.2, we train our model using the source dataset and directly evaluate it on the target dataset. In this task, we choose M4 as the source dataset and M3 as the target dataset. It is important to note that the structure of the M3 and M4 datasets differs from typical datasets used for long-term forecasting. They consist of multiple univariate time sequences of different lengths. The final part of each sequence serves as the test set, while the preceding part is used for training.
In zero-shot scenarios, we use a sequence of length $F$ preceding and consecutive to the test set as input, referred to as the lookback window, where $F$ is the forecast length of each subset. During in-context forecasting, we concatenate the first $2F$ time points that belong to the same sequence with the lookback window as input. We aim to enhance prediction performance by incorporating more contextual information. Too short sequences ($<4F$) are discarded to prevent overlap between the prompt and the lookback window. For a fair comparison, both zero-shot and in-context forecasting performance are reported on the same remaining sequences. Figure 11 provides showcases of zero-shot and in-context forecasting.
@@
  • P.1 v.s. P.4 indicates that the selected prompt is not suitable, since the prompt does not come from the earlier observations of the same series to be predicted. Although the context window becomes larger, the averaged performance will still deteriorate.
  • P.2 v.s. P.3 indicates that in most cases, selecting the last $2F$ time series can provide better contextual information, but not always.
This highlights the prompt engineering for in-context forecasting. An intuitive suggestion is to utilize consecutive, inter-periodic, and multiple prompts, which is promising and leaves future work.

### v4

@@ 4.1 Time Series Forecasting
For short-term forecasting, we follow the well-acknowledged TimesNet [42], which assesses the fundamental ability of forecasters in modeling temporal variations. For long-term forecasting, we establish a novel one-for-all benchmark: a single forecaster is trained on one dataset and subsequently utilized for all prediction lengths. We highlight that this approach evaluates the basic versatility as foundation models of time series, which aims to break the prevailing practice of extensive training across diverse real-world scenarios. To be specific, we evaluate all methods by rolling forecasting: a model is trained with predetermined input/output lengths, and the predicted values are integrated as part of the input in subsequent iterations until reaching the desired forecast length. Therefore, the key to success in this task lies in mitigating multi-step error accumulation. Still, the conventional one-for-one approach that trains forecasters respectively on each length is also provided in Table 12.
@@ Results
By diving into the proposed one-for-all and the traditional one-for-one benchmarks in Table 3 and 12, it is notable that prevalent deep forecasters, such as Transformer-based forecasters and DLinear, can achieve competitive and even better results under rolling forecasting. Nevertheless, the performance of non-autoregressive LLM4TS methods can degenerate a lot without respective training. Therefore, it highlights our persistent utilization of autoregression and thorough leveraging of inherent token transitions of LLMs, thereby mitigating error accumulation during multi-step rolling forecasting.
@@ 4.3 In-Context Forecasting
We adopt the aforementioned M4 $\to$ M3 scenario. Since the samples of the M3 dataset are univariate time series with different lengths, we always predict the last $F$ time points of each sample during inference. We set the lookback length $L=F$, and thus the length of time series prompts is $2F$. We set the number of prompts $m=1$. Therefore, we initially train an LLM-based forecaster on the source M4 dataset with the context length of $3F$. We adopt an intuitive strategy to select the prompt: uniformly adopting the first $2F$ time points of the time series as the corresponding prompt. Supposing the lookback series starts after time $t\ (\geq F)$, in-context forecasting is formulated as:
$f:(\{x_{1:2F}\},x_{t+1:t+F},\mathbf{a}_{t+1:t+2F})\mapsto\hat{x}_{t+F+1,t+2F}.$ (12)
To prevent data leakage of future information, too short samples are discarded to prevent the overlap between the prompt and the future prediction. The implementation details of in-context forecasting are provided in Appendix D.6. We further investigate different prompt retrieval strategies. Insightful results are provided to reveal the influence of using time series prompts for interactive prediction and take-away instructions of prompt engineering in the time series modality.
@@ 4.4 Method Analysis
Ablation study
Recent research has raised doubts about the validity of previous LLM4TS methods [35], which predominantly adopt non-autoregression, that is, treating the LLM as a BERT-style pre-trained backbone and utilize a globally flattening projector on all lookback tokens. Here, we provide a thorough ablation study to examine our proposed AutoTimes in Table 6. The results underscore that our method maintains the consistency of the decoder-only architecture and autoregressive inference, effectively leveraging LLMs and addressing the concerns regarding performance improvement and adaptation cost. Further, we provide a comparison by substituting our token-wise segment projection (consistent with LLMs) with the flatten linear head [26] (common in non-autoregressive forecasters). Results of Table 21 in the Appendix reveal that the performance of non-autoregressive generation is consistently inferior to that of our autoregressive AutoTimes approach.
Table 6: We follow the protocol of LLM4TS ablation studies [35] to verify whether the LLM is truly useful in our AutoTimes: (1) w/o LLM replaces the language model entirely and passing input tokens directly to the last layer; (2) LLM2Attn replaces the language model with a single multi-head attention layer; (3) LLM2Trsf replaces the language model with a single transformer block.
Dataset ETTh1 ECL
Type
AutoTimes
w/o LLM
LLM2Attn
LLM2Trsf
AutoTimes
w/o LLM
LLM2Attn
LLM2Trsf
Metric
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
Pred-$96$
0.360
0.400
0.365
0.399
0.383
0.404
0.377
0.401
0.129
0.225
0.171
0.263
0.156
0.255
0.162
0.263
Pred-$192$
0.388
0.419
0.405
0.425
0.414
0.422
0.406
0.420
0.147
0.241
0.192
0.282
0.178
0.276
0.189
0.287
Pred-$336$
0.401
0.429
0.429
0.441
0.431
0.432
0.421
0.431
0.162
0.258
0.216
0.304
0.198
0.295
0.216
0.309
Pred-$720$
0.406
0.440
0.450
0.468
0.456
0.454
0.449
0.452
0.199
0.288
0.264
0.342
0.230
0.320
0.258
0.340
@@
Table 7: Full long-term forecasting results of AutoTimes and AutoTimes equipped with LoRA [14].
Dataset
@@
Pred-$96$ AutoTimes 0.360 0.397 0.140 0.236 0.158 0.208 0.369 0.257 0.179 0.220
+ LoRA
0.357 0.397 0.130 0.225 0.151 0.201 0.360 0.256 0.176 0.219
Pred-$192$ AutoTimes 0.391 0.419 0.159 0.253 0.207 0.254 0.394 0.268 0.198 0.236
+ LoRA
0.391 0.420 0.149 0.242 0.197 0.244 0.383 0.267 0.195 0.235
Pred-$336$ AutoTimes 0.408 0.432 0.177 0.270 0.262 0.298 0.413 0.278 0.213 0.252
+ LoRA
0.409 0.433 0.164 0.259 0.251 0.287 0.401 0.277 0.208 0.249
Pred-$720$ AutoTimes 0.429 0.452 0.216 0.303 0.342 0.353 0.449 0.299 0.239 0.277
+ LoRA
0.426 0.451 0.202 0.293 0.326 0.339 0.440 0.300 0.225 0.268
@@ 5 Conclusion
Acknowledgments
This work was supported by the Ministry of Industry and Information Technology of China.
@@ Appendix A Dataset Descriptions
Table 8: Detailed dataset descriptions. Dim denotes the variate number. Dataset Size denotes the total number of time points in (Train, Validation, Test) splits respectively. Forecast Length denotes the future time points to be predicted. Frequency denotes the sampling interval of time points.
@@ Appendix B Implementation Details
AutoTimes processes timestamps in textual form rather than numerical encoding, potentially enabling to handle other textual data such as news or logs. We utilize LLM to obtain embedding for the special token <EOS> to capture embedding for the entire sentence. Pseudo-code for this process is depicted in Algorithm 1. It is worth noting that in the context of multivariate time series forecasting, timestamps are shared across variates. Thus, timestamps can implicitly express relationships between variates even with channel independence. Further, assuming there are $C$ variates since the number of timestamps is $\frac{1}{C}$ of the total time point count, these embeddings can be efficiently pre-computed by large language models.
After obtaining embedding for the timestamps, we repurpose LLM for time series forecasting using Algorithm 2. At this stage, only the parameters of $\operatorname{SegmentEmbedding}$ and $\operatorname{SegmentProjection}$ are updated, while the parameters of LLMs remain entirely frozen. During inference, AutoTimes utilizes the last token generated as its prediction and then employs this output to create subsequent predictions autoregressively. This approach enables AutoTimes to predict sequences of variable lengths with just one model dynamically. Such capability is crucial in real-world application scenarios. The pseudo-code in Algorithm 3-4 illustrates this process.
@@
2:$\mathbf{s}_{i}=\{x_{(i-1)S+1},\dots,x_{iS}\}$triangleright Model dimension of the LLM $D$
@@
1:Input time series $\{x_{1},\dots,x_{N_{1}\times S}\}$; textual embeddings $\{\mathbf{TE}_{1},\dots,\mathbf{TE}_{N_{2}}\}$triangleright Forecast token number $N_{2}-N_{1}$
@@ D.1 Time Series Forecasting
We compare the performance of AutoTimes with state-of-the-art LLM4TS methods and well-acknowledged deep forecasters. Table 11 shows detailed short-term forecast results on M4. Table 10 presents results of the one-for-all forecasting benchmark across ETTh1, ECL, Traffic, Weather, and Solar-Energy datasets. We evaluate all methods by rolling forecasting: each model is trained with input length $672$ and output length $96$, and the predicted values are integrated as part of the input in the next iteration until reaching the desired forecast horizon.
@@
Additionally, we evaluate AutoTimes along with other baseline models on recent benchmarks [24]. Results are presented in Table 15. We also look forward to evaluating on more diverse benchmarks in the future.
@@
Method AutoTimes
TimeLLM [15]
UniTime [21]
FPT [49] iTrans. [22] DLinear [45]
PatchTST [26]
TimesNet [42]
@@
Table 12: Long-term forecasting results of one-for-one: AutoTimes trains one LLM-based forecaster to handle all prediction lengths by autoregression, whereas other models are trained respectively on each prediction length. AutoTimes adapts LLMs with the context length $C=672$. The lookback length is set as $L=672$ in others. All results are averaged. Full results is provided in Table 14.
@@
Table 15: Forecasting results on additional benchmark datasets [24] ($672$-pred-$96$).
Models AutoTimes PatchTST iTransformer DLinear
Metric MSE MAE MSE MAE MSE MAE MSE MAE
Australian Electricity 0.150 0.228 0.163 0.242 0.153 0.233 0.167 0.250
Bdg-2 Panther 0.537 0.458 0.565 0.476 0.546 0.462 0.581 0.499
Oikolab Weather 0.603 0.577 0.635 0.603 0.630 0.591 0.663 0.611
@@ D.2 Zero-Shot Forecasting
Following the zero-shot forecasting of FPT [15], each experiment comprises the source and target datasets. We train a model on the source dataset and apply the model on the target dataset for predictions directly.
@@ D.3 Method Generality
Mainstream LLMs predominantly adopt the decoder-only architecture, and AutoTimes can utilize any decoder-only LLM. We conduct experiments on various types and sizes of LLMs, including GPT-2 [31], multiple sizes of OPT [46], and LLaMA [36]. Detailed configurations and results are shown in Table 17 and 18, demonstrating a general trend where performance improves as the model size increases, consistent with the scaling law [16].
Table 17: Detailed method configurations of AutoTimes for alternative language models.
Base LLM GPT-2 (124M) OPT-350M OPT-1.3B OPT-2.7B OPT-6.7B LLaMA-7B
Hidden Dim. 768 1024 2048 2560 4096 4096
Embedding 2-layer MLP 2-layer MLP 2-layer MLP 2-layer MLP 2-layer MLP Linear
Trainable Param. (M) 0.44 0.58 1.10 1.36 2.15 0.79
@@ D.4 Variable Lookback Length
In the conventional forecasting paradigm, deep forecasters are trained respectively on lookback/forecast lengths, limiting their applicability to a single lookback length. In contrast, LLMs have the versatility to handle various input lengths. This capability is derived from Rotary Position Embedding  [33] and the next toke prediction objective, where LLMs are trained with token-wise supervision in Equation 8, that is, the generated token at each position is supervised. While non-autoregressive LLM4TS methods are typically constrained to a fixed lookback setting, AutoTimes with a consistent training objective has the flexibility to deal with different lookback lengths. We present the results in Figure 8, where we adapt the LLM by AutoTimes with the context length of $C=672$, and evaluate the performance with different lookback lengths, which demonstrates the inherited versatility of LLM-based forecasters. Moreover, the performance is generally improving with increased available lookback observations, leading to an averaged $9.3\%$ promotion from $384$ to $672$. By contrast, several works have observed that the performance of respectively trained deep forecasters does not necessarily improve with the increasing of lookback length [22, 26, 45].
@@ D.6 In-Context Forecasting
For in-context forecasting, similar to zero-shot forecasting in Appendix D.2, we train our model using the source dataset and directly evaluate it on the target dataset. In this task, we first choose M4 as the source dataset and M3 as the target dataset. It is important to note that the structure of the M3 and M4 datasets differs from typical datasets used for long-term forecasting. They consist of multiple univariate time sequences of different lengths. The final part of each sequence serves as the test set, while the preceding part is used for training.
Implementation
In zero-shot scenarios, we use a sequence of length $F$ preceding and consecutive to the test set as input, referred to as the lookback window, where $F$ is the forecast length of each subset. During in-context forecasting, we concatenate the first $2F$ time points that belong to the same sequence with the lookback window as input. We aim to enhance prediction performance by incorporating more contextual information. Too short sequences ($\leq 4F$) are discarded to prevent overlap between the prompt and the lookback window. For a fair comparison, both zero-shot and in-context forecasting performance are reported on the same remaining sequences. Figure 12 provides showcases of zero-shot and in-context forecasting.
@@
  • P.1 v.s. P.4 indicates that the selected prompt is not suitable, since the prompt does not come from the earlier observations of the same series to be predicted. Although the context window becomes larger, the averaged performance will deteriorate because of irrelevant prompts.
  • P.2 and P.3 indicates that in most cases, selecting the relevant $2F$ time series from the same series can provide better contextual information.
This highlights the prompt engineering for in-context forecasting. An intuitive suggestion is to utilize consecutive, inter-periodic, and multiple prompts. To verify this idea, we analyze the periodic effect of time series prompts.
@@
In previous experiments, we adopt M3 and M4 datasets, which are consistent with the zero-shot experiment of FPT [49], to present the promotion of our in-context paradigm. To provide more rigorous conclusions, we extend the evaluation to widely recognized datasets. Details of the experiment are as follows: By using a trained model checkpoint on a source domain (Traffic), we conduct forecasting without gradient update on target ETT datasets. We evaluate the pred-$96$ performance on the last variate (OT). For the zero-shot scenario, the input is length-$288$ lookback series. For in-context forecasting, the input is (length-$384$ series prompt + length-$288$ lookback series). Considering the dataset periodicity, the prompt is uniformly selected as the Ahead-24 (one-day-ahead) series of the original lookback series. To eliminate the performance boost that comes from extending the input length, we also provide the results of length-$672$ lookback series in the zero-shot scenario. Moreover, we further delve into the effect of different strategies to select time series prompts:
  • Ahead-Period: The prompt is uniformly selected as the Ahead-$24$ series of the original lookback series where $24$ is one of the periods (daily period) of ETT.
  • Ahead-Random: The prompt is randomly selected as the previous series of the original series.
  • Fixed Prompt: The prompt is fixed as the first $384$ time points in the variate-OT.
  • Other Variate: The prompt is uniformly selected as Ahead-$24$ series, but comes from other variates.
Table 20: Strategies to select time series prompts based on periodicity for in-context forecasting.
Context for prediction
ETTh1-OT
ETTh2-OT
ETTm1-OT
ETTm2-OT
Average Err.
P.0: Zero-Shot (Input-$288$))
0.0673 0.1637 0.0424 0.1669 0.1101
P.1: Zero-Shot (Input-$672$)
0.0657 0.1538 0.0415 0.1701 0.1078
P.2: Ahead-Period (Input-$672$)
0.0645 0.1513 0.0399 0.1629 0.1047
P.3: Ahead-Random (Input-$672$)
0.0666 0.1621 0.0407 0.1719 0.1103
P.4: Fixed Prompt (Input-$672$)
0.0769 0.1859 0.0512 0.2104 0.1311
P.5: Other-Variates (Input-$672$)
0.1263 0.1780 0.0852 0.2297 0.1548
Results in Table 20 demonstrate the effectiveness of using suitable time series prompts and highlight the influence of prompt engineering. Using inter-period prompts can outperform simply extending lookback window.
The benefit of the proposed in-context forecasting is to extend the input context of time series forecasting beyond a continuous lookback window. Since the essence of prompts is to incorporate useful domain-specific knowledge, here is one use case of in-context forecasting: Considering predicting the weather of one day, one approach is to extend the lookback length from days to weekends. However, it can also introduce noisy information since non-stationary meteorological conditions can change with seasons. Another practical way is to consider how the weather changes on the same day in the last year (or years). Although the input is not continuous, the input context becomes more relevant based on prior knowledge about the periodicity (yearly). Therefore, in-context forecasting makes prior knowledge incorporatable and gets performance promotion.
D.7 Ablation Study
In addition to the ablation study of whether LLMs are useful in AutoTimes (Table 6), we further delve into the main difference between our method and previous LLM4TS approach and provide a comprehensive ablation study. The results presented in Table 21 demonstrate that the performance of non-autoregression projection is consistently inferior to that of our autoregressive AutoTimes approach.
Table 21: Ablation study of the autoregression. FlattenHead replaces the segment-wise projection of AutoTimes by flatten and linear head [26], which is prevalent in non-autoregressive forecasters.
Dataset ETTh1 ECL Weather Traffic
Type
AutoTimes
FlattenHead
AutoTimes
FlattenHead
AutoTimes
FlattenHead
AutoTimes
FlattenHead
Metric
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
Pred-$96$
0.360
0.400
0.385
0.420
0.129
0.225
0.142
0.247
0.153
0.203
0.155
0.209
0.343
0.248
0.367
0.261
Pred-$192$
0.388
0.419
0.445
0.463
0.147
0.241
0.157
0.259
0.201
0.250
0.202
0.251
0.362
0.257
0.391
0.282
Pred-$336$
0.401
0.429
0.463
0.475
0.162
0.258
0.201
0.311
0.256
0.293
0.257
0.286
0.379
0.266
0.404
0.287
Pred-$720$
0.406
0.440
0.574
0.542
0.199
0.288
0.232
0.331
0.331
0.345
0.333
0.261
0.413
0.284
0.432
0.294